OpenStudy (anonymous):

Hannah has a jar filled with 150 coins made up of quarters and nickels. The total value of the coins is $28.30. Let x represent the number of quarters in the jar. Which equation can be used to find the number of quarters?

OpenStudy (anonymous):

A. 0.25x + 0.05(x – 150) = 28.30 B. 0.05x + 0.25(x – 150) = 28.30 C. 0.05x + 0.25(150 – x) = 28.30 D. 0.25x + 0.05(150 – x) = 28.30

OpenStudy (anonymous):

I see you're typing and it feels like the Lord is answering my prayers.

OpenStudy (anonymous):

solve the system of equations